Shift report 20010926 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Sep 26 2001 - 07:48:51 EDT

  • Next message: Apache: "Shift report 20010926 08:00-16:00"

    ------------------------------------------------------------
    Date: Wed Sep 26 07:48:51 2001, Shift: 00:00-08:00
    Supervisor: JN, with: 
    Log book#9 pages 48 - 49
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam from 00:00-05:35
        No Beam 05:35-07:30
        Filling RHIC began at 07:30
    2.  Data Runs 4953, 4954
        Bad Pedestal Runs 4955, 4956 (incorrect scaledowns)
        (Was also turning down T3, T4, and T5 Anode voltages 
         during 4956 - is this bad?!?)
        Pedestal Run 4957
        TPC Pedestal Run 4958
        DVM Run 4959
    3.  Still having "unexecpted reply from mainframe" for
        BBL, Tile, TFCAL, TPM1, TPM2, T1 and T2.
        Gating A,B,C still off on TPM1, TPM2, T1 and T2.
        LEDs still not blinking on TPM1, TPM2, T1 and T2.
    
    And on a personal note, this is my last graveyard shift 
    for this stint - thank you very much.  And with that 
    note: Elvis leaves the building.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Sep 26 2001 - 07:49:19 EDT